OpenStudy (anonymous):

Given f(x) = –3x – 7, what is f(–4)? A. 5 B. –14 C. –1 D. 40

OpenStudy (anonymous):

plug the value of -4 into the variables (x) so -3(-4) - 7 = f(x) f(x) = 12 - 7 f(x) = 5

OpenStudy (solomonzelman):

yes, but don't say that this is the value of f(x), say f(-4)